# An electron-hadron collider at the high-luminosity LHC

### v2

@@ 3 Physics case
In the phase-one LHeC, the centre-of-mass energy for electron-proton collisions $\sqrt{s_{ep}}=\sqrt{4E_{e}E_{p}}=0.75$ TeV, in place of 1.2 TeV for the 50 GeV electron beam at the final LHeC. This means that the kinematic reach on the $(x,Q^{2})$ plane will already be greatly extended with respect to the data taken at HERA at $\sqrt{s}=0.32$ TeV, as $Q^{2}_{\rm max}\simeq s_{ep}$ and $x_{\rm min}\propto 1/s_{ep}$. As discussed in the LHeC CDR [2], a relatively low integrated luminosity of $50~{\rm fb}^{-1}$ is sufficient to greatly improve the constraints on the parton density functions, particularly at low $x$. The longitudinal electron polarisations of both signs will be instrumental in decomposing the parton distributions, and in particular, in determining the total contribution of sea quarks and their possible charge asymmetry, as well as for measuring the quark density ratio $u/d$. One should stress that it also includes studies of the nuclear parton distributions. Therefore, already after two years of running the phase one LHeC it will be possible to provide valuable feedback to hadron-hadron experiments at the HL-LHC, which will result in significant improvements in a number of precision measurements [4]. Moreover, with entering into a new kinematic region of very low $x$ the novel QCD phenomena are expected to be observed – new parton dynamics and non-linear parton density effects.
@@
$\sigma$($e^{-}p\to e^{-}\,t\bar{t}~+$ X) [fb] 26 28 0.93 0.99
$\sigma$($e^{-}p\to\nu_{l}\,\bar{t}~+$ X) [pb] 1.7 3.1 0.28 0.51
To illustrate the scientific potential of the research in electroweak and Higgs physics at the phase-one LHeC, in Tab. 3 we compare four generator-level cross-sections of processes of interest for un-polarised and longitudinally polarised electron beams to those for the final LHeC. All the values cited here are obtained using the version 3.5.8 of the MadGraph5_aMC@NLO [14] event generator, using the CTEQ6 [15] set of parton density functions evaluated at leading order. The Higgs boson case remains compelling at 0.75 TeV (especially for the charged current events), as well as the single-top-quark one, not only for mastering pioneering electron-proton measurement techniques but also for making measurements of some specific channels. For example, assuming the Run5 total integrated luminosity of $250~{\rm fb}^{-1}$ at 0.75 TeV, and a 80-80% electron longitudinal polarisation, we obtain the following uncertainties by scaling up the errors obtained for the 1.2 TeV LHeC [2] using the numbers of produced events: in the nomenclature of the $\kappa$ framework of an effective field theory (EFT) extension in the Higgs sector [16], $\delta\kappa_{W}\approx 2.5\%$ for the Higgs boson coupling to $W$ bosons, and $\delta{\cal R}e(f^{2}_{L})\approx 0.2$, for the anomalous $Wtb$ coupling in dimension-6 EFT extensions [17] – both errors comparable to the expected final uncertainties for the corresponding measurements at the HL-LHC using $3~{\rm ab}^{-1}$ of the $pp$ data.
In Fig. 7 we show the differential cross-sections in rapidity for the Higgs boson and a single top quark produced at the phase-one LHeC as well as the final one – for the latter the cross-sections peak at about $Y=2$ requiring good coverage in the forward part of the hadron hemisphere. For the main decay products of these two exemplary processes of interest, a pseudo-rapidity coverage of up to 4 units, both in the tracking and calorimetry components of this hemisphere, would ensure a proper detection efficiency. The big advantage of using the highly polarised electron beam is clearly visible.
@@
Experimental conditions at the LHeC will be particularly favourable for studies of the exclusive production via photon-photon fusion, the event pile-up will be negligible, highly improving the efficiency and purity of exclusive event selection, and at the same time some sources of overwhelming backgrounds, as due to the Drell-Yan process for example, will be absent. Moreover, no event suppression will occur due to hadronic re-scattering and the associated significant theoretical and experimental uncertainties will not intervene. Finally, if we consider the usage of data streaming (or the absence of triggering) in acquisition, the detection and reconstruction of a much wider range of final states will be allowed. As a result, despite the lower average $\gamma\gamma$ centre-of-mass energy than that at the HL-LHC, a significantly larger statistics of detected photon-photon interactions is expected already at the phase-one LHeC [18].
@@
For example, using the CepGen v1.2.2 [19] implementation of the LPAIR [20] matrix element-level event generator, we estimate a large cross-section for the fully exclusive two-photon production of tau-lepton pairs: about 40 pb at the phase-one LHeC, for pair invariant masses above 10 GeV. This will lead to very large event statistics within the acceptance of the central detector and as a consequence to excellent sensitivity to the $\tau$ anomalous magnetic moment $a_{\tau}$. With an integrated $ep$ luminosity of 250 fb-1, the expected sensitivity is an order of magnitude better than that achieved at the LHC [21]. As a result, for the first time the experimental uncertainties will be better than the higher-order corrections to the $\tau$ magnetic moment in the Standard Model. Similarly, large improvements will be achieved in constraining the electric dipole moment of tau-leptons.
@@
Of course, there are many more searches for the BSM physics to be performed at the LHeC as, for example, unique searches for supersymmetric particles and studies of anomalous quartic gauge couplings [18]. Moreover, signatures such as displaced vertices or short tracks may emerge, which can be efficiently identified at the LHeC due to the low level of hadronic background and event pile-up with respect to the LHC, and as such offering a considerable window for discovery. Notable examples involve heavy neutrinos and axion-like particles, inspired by the QCD axion [2, 4].
@@
Last but not least, already at the phase-one LHeC, also one of the key questions in nuclear physics can be addressed: how the nucleon structure is modified when immersed in a nuclear medium. For example, by a precise measurement and complete unfolding of the distributions of all parton species in a single nucleus comparable to that in the proton, the LHeC will revolutionise our understanding of nuclear structure in crucial and unique kinematic regions [2, 4]. In addition, it will allow to clarify the dynamics for particle production in nuclear collisions in the region of relevance for the LHC. Finally, it should be stressed that the phase-one LHeC provides an important opportunity to perform $eA$ and $AA$ measurements using the same detector which provides the unique advantage for cross-calibration of performance and physics.
@@
The overlap in the physics programmes of the LHeC and the Electron-Ion Collider (EIC) at Brookhaven is small, as the two projects will produce very important but also very complementary scientific outputs. The maximal centre-of-mass energy of electron-proton collisions at the EIC is 0.14 TeV therefore the electroweak and Higgs physics play there only marginal role, in stark contrast to the LHeC. Moreover, the significantly higher centre-of-mass energy at the LHeC corresponds to a much wider coverage on the $(x,Q^{2})$ plane, as explained above, allowing for unique studies of non-linear effects in partonic interactions and for significantly improved predictions of low-$x$ processes in the $pp$ collisions at the LHC. On the other hand, both beams can be polarised at the EIC and that will provide means for new insights into the origin of spin in hadrons, in particular. Last but not least, there is a great synergy between the two projects in the detector developments and of the associated electronics.
@@ 4 General detector requirements
In 2022, an extension of the ALICE experiment beyond Run4 at the HL-LHC was proposed in a letter of intent [22]. In that document, two configurations of magnetic field were considered for the new ALICE 3 detector at the IP2 – produced either by the cost effective 7.5 m long solenoid, or by the high-performance combination of a 2.5 m solenoid at the centre with two 2 m long dipoles on its each side. As discussed above, to allow for the optimal electron-hadron collisions, the dipole field is necessary for the separation of the electron beam, so the latter approach is desirable in this context.
The most recent ALICE 3 design includes a tracking detector with excellent performance that covers a wide range of rapidities of $|\eta|<4$ [23], which is well compatible with the needs for electron-hadron measurements at the LHeC with 20 GeV electrons. However, in contrast, the present design lacks a hadronic calorimeter (HCAL) while a performant one is obligatory for performing proper measurements of the DIS charged current events and reconstruction of the decays of top quark, Higgs, $W$ and $Z$ bosons. Moreover, an extension of the electromagnetic calorimeter is desirable in the electron forward direction to ensure a good reconstruction of DIS neutral current events; in addition, important studies of exclusive production at the LHeC would largely benefit from specialised near-beam detectors in very forward electron and proton directions.
@@ 5 Conclusions
Acknowledgements
We thank T. M. P. von Witzleben for her contribution on the study of the beta-beat in the full HL-LHC lattice. L. Forthomme and K. Piotrzkowski appreciate the financial support of the Polish National Agency for Academic Exchange (NAWA) under grant number BPN/PPO/2021/1/00011. L. Forthomme has also been supported by the NCN grants 2022/01/1/ST2/00022 and 2023/49/B/ST2/03273, and by the AGH IDUB and the CERN COAS programmes.

### v3

@@ 3 Physics case
In the phase-one LHeC, the centre-of-mass energy for electron-proton collisions $\sqrt{s_{ep}}=\sqrt{4E_{e}E_{p}}=0.75$ TeV, in place of 1.2 TeV for the 50 GeV electron beam at the final LHeC. This means that the kinematic reach on the $(x,Q^{2})$ plane will already be greatly extended with respect to the data taken at HERA at $\sqrt{s}=0.32$ TeV [Klein:2008di], as $Q^{2}_{\rm max}\simeq s_{ep}$ and $x_{\rm min}\propto 1/s_{ep}$. As discussed in the LHeC CDR [LHeC:2020van], a relatively low integrated luminosity of $50~{\rm fb}^{-1}$ is sufficient to greatly improve the constraints on the parton density functions, particularly at low $x$. The longitudinal electron polarisations of both signs will be instrumental in decomposing the parton distributions, and in particular, in determining the total contribution of sea quarks and their possible charge asymmetry, as well as for measuring the quark density ratio $u/d$. One should stress that it also includes studies of the nuclear parton distributions. Therefore, already after two years of running the phase one LHeC it will be possible to provide valuable feedback to hadron-hadron experiments at the HL-LHC, which will result in significant improvements in a number of precision measurements [Ahmadova:2025vzd]. Moreover, with entering into a new kinematic region of very low $x$ the novel QCD phenomena are expected to be observed – new parton dynamics and non-linear parton density effects.
@@
$\sigma$($e^{-}p\to e^{-}\,t\bar{t}~+$ X) [pb] 2.1 2.2 0.45 0.48
$\sigma$($e^{-}p\to\nu_{l}\,\bar{t}~+$ X) [pb] 5.8 10.7 0.74 1.34
To illustrate the scientific potential of the research in electroweak and Higgs physics at the phase-one LHeC, in Tab. 3 we compare four generator-level cross-sections of processes of interest for un-polarised and longitudinally polarised electron beams to those for the final LHeC. All the values cited here are obtained using the version 3.5.8 of the MadGraph5_aMC@@NLO [Alwall:2014hca] event generator, using the CTEQ6 [Pumplin:2002vw] set of parton density functions evaluated at leading order. The Higgs boson case remains compelling at 0.75 TeV (especially for the charged current events), as well as the single-top-quark one, not only for mastering pioneering electron-proton measurement techniques but also for making measurements of some specific channels. For example, assuming the Run5 total integrated luminosity of $250~{\rm fb}^{-1}$ at 0.75 TeV, and a 80-80% electron longitudinal polarisation, we obtain the following uncertainties by scaling up the errors obtained for the 1.2 TeV LHeC [LHeC:2020van] using the numbers of produced events: in the nomenclature of the $\kappa$ framework of an effective field theory (EFT) extension in the Higgs sector [Ghezzi:2015vva], $\delta\kappa_{W}\approx 2.5\%$ for the Higgs boson coupling to $W$ bosons, and $\delta{\cal R}e(f^{2}_{L})\approx 0.2$, for the anomalous $Wtb$ coupling in dimension-6 EFT extensions [Chen:2005vr] – both errors comparable to the expected final uncertainties for the corresponding measurements at the HL-LHC using $3~{\rm ab}^{-1}$ of the $pp$ data.
In Fig. 7 we show the differential cross-sections in rapidity for the Higgs boson and a single top quark produced at the phase-one LHeC as well as the final one – for the latter the cross-sections peak at about $Y=2$ requiring good coverage in the forward part of the hadron hemisphere. For the main decay products of these two exemplary processes of interest, a pseudo-rapidity coverage of up to 4 units, both in the tracking and calorimetry components of this hemisphere, would ensure a proper detection efficiency [^1] The large beam energy asymmetry at the LHeC is well compensated by such a wide $\eta$ coverage, even for the 20 GeV beam – the electron-parton interactions are kinematically equivalent to the parton-parton interactions with the ”electron” $x=20/7000\simeq 0.003$, which is quite typical at the LHC.. The big advantage of using the highly polarised electron beam is clearly visible.
@@
Experimental conditions at the LHeC will be particularly favourable for studies of the exclusive production via photon-photon fusion, the event pile-up will be negligible, highly improving the efficiency and purity of exclusive event selection, and at the same time some sources of overwhelming backgrounds, as due to the Drell-Yan process for example, will be absent. Moreover, no event suppression will occur due to hadronic re-scattering and the associated significant theoretical and experimental uncertainties will not intervene. Finally, if we consider the usage of data streaming (or the absence of triggering) in acquisition, the detection and reconstruction of a much wider range of final states will be allowed. As a result, despite the lower average $\gamma\gamma$ centre-of-mass energy than that at the HL-LHC, a significantly larger statistics of detected photon-photon interactions is expected already at the phase-one LHeC [Forthomme:2024qqi].
@@
For example, using the CepGen v1.2.2 [Forthomme:2018ecc] implementation of the LPAIR [Vermaseren:1982cz] matrix element-level event generator, we estimate a large cross-section for the fully exclusive two-photon production of tau-lepton pairs: about 40 pb at the phase-one LHeC, for pair invariant masses above 10 GeV. This will lead to very large event statistics within the acceptance of the central detector and as a consequence to excellent sensitivity to the $\tau$ anomalous magnetic moment $a_{\tau}$. With an integrated $ep$ luminosity of 250 fb-1, the expected sensitivity is an order of magnitude better than that achieved at the LHC [CMS:2024qjo]. As a result, for the first time the experimental uncertainties will be better than the higher-order corrections to the $\tau$ magnetic moment in the Standard Model. Similarly, large improvements will be achieved in constraining the electric dipole moment of tau-leptons.
@@
Of course, there are many more searches for the BSM physics to be performed at the LHeC as, for example, unique searches for supersymmetric particles and studies of anomalous quartic gauge couplings [Forthomme:2024qqi]. Moreover, signatures such as displaced vertices or short tracks may emerge, which can be efficiently identified at the LHeC due to the low level of hadronic background and event pile-up with respect to the LHC, and as such offering a considerable window for discovery. Notable examples involve heavy neutrinos and axion-like particles, inspired by the QCD axion [LHeC:2020van, Ahmadova:2025vzd].
@@
Last but not least, already at the phase-one LHeC, also one of the key questions in nuclear physics can be addressed: how the nucleon structure is modified when immersed in a nuclear medium. For example, by a precise measurement and complete unfolding of the distributions of all parton species in a single nucleus comparable to that in the proton, the LHeC will revolutionise our understanding of nuclear structure in crucial and unique kinematic regions [LHeC:2020van, Ahmadova:2025vzd]. In addition, it will allow to clarify the dynamics for particle production in nuclear collisions in the region of relevance for the LHC. Finally, it should be stressed that the phase-one LHeC provides an important opportunity to perform $eA$ and $AA$ measurements using the same detector which provides the unique advantage for cross-calibration of performance and physics.
@@
The overlap in the physics programmes of the LHeC and the Electron-Ion Collider (EIC) at Brookhaven is small, as the two projects will produce very important but also very complementary scientific outputs. The maximal centre-of-mass energy of electron-proton collisions at the EIC is 0.14 TeV therefore the electroweak and Higgs physics play there only marginal role [AbdulKhalek:2021gbh], in stark contrast to the LHeC. Moreover, the significantly higher centre-of-mass energy at the LHeC corresponds to a much wider coverage on the $(x,Q^{2})$ plane, as explained above, allowing for unique studies of non-linear effects in partonic interactions and for significantly improved predictions of low-$x$ processes in the $pp$ collisions at the LHC. On the other hand, both beams can be polarised at the EIC and that will provide means for new insights into the origin of spin in hadrons, in particular. Last but not least, there is a great synergy between the two projects in the detector developments and of the associated electronics.
@@ 4 General detector requirements
In 2022, an extension of the ALICE experiment beyond Run4 at the HL-LHC was proposed in a letter of intent [ALICE:2022wwr]. In that document, two configurations of magnetic field were considered for the new ALICE 3 detector at the IP2 – produced either by the cost effective 7.5 m long solenoid, or by the high-performance combination of a 2.5 m solenoid at the centre with two 2 m long dipoles on its each side. As discussed above, to allow for the optimal electron-hadron collisions, the dipole field is necessary for the separation of the electron beam, so the latter approach is desirable in this context.
The most recent ALICE 3 design includes a tracking detector with excellent performance that covers a wide range of rapidities of $|\eta|<4$ [Dainese:2925455], which is well compatible with the needs for electron-hadron measurements at the LHeC with 20 GeV electrons. However, in contrast, the present design lacks a hadronic calorimeter (HCAL) while a performant one is obligatory for performing proper measurements of the DIS charged current events and reconstruction of the decays of top quark, Higgs, $W$ and $Z$ bosons. Moreover, an extension of the electromagnetic calorimeter is desirable in the electron forward direction to ensure a good reconstruction of DIS neutral current events; in addition, important studies of exclusive production at the LHeC would largely benefit from specialised near-beam detectors in very forward electron and proton directions.
@@ 5 Conclusions
[…]
No new data were created or analysed in this study.
[…]
We thank T. M. P. von Witzleben for her contributions to the study of the LHeC optics.
[…]
L. Forthomme and K. Piotrzkowski appreciate the financial support of the Polish National Agency for Academic Exchange (NAWA) under grant number BPN/PPO/2021/1/00011. L. Forthomme has also been supported by the NCN grants 2022/01/1/ST2/00022 and 2023/49/B/ST2/03273, and by the AGH IDUB and the CERN COAS programmes.